Project

General

Profile

6/1/2015 Notes

DocDB Event

Attendance

Rob, Eric, Jan, Serdar, Leon, Andrey, Cristiana, Enhao and Martin (taking notes)

Cooling

  • Leon explains that there was a problem with the relay for the compressor, so every once in a while it would cut out.
  • This relay has now been fixed, so we will increase the number of DAQ nodes the next time a run is restarted.

Data Quality

  • You get different rates for strips in the horizontal and vertical direction.
  • When the beam comes back on, the rates then get smeared.

DDT Deployment

  • The slow monopole trigger will be updated and deployed this week.
  • If possible, Yuri will update the Up-Mu trigger to move the art association call outside of the loop, so we can pull it into this update.
  • There is no new DAQ release yet, so we can unfortunately not pull in the Supernova and Moon triggers.
  • Once the DAQ group cuts a new FD release, I will gladly deploy again.